### v3

# Bearing-Distance Based Flocking with Zone-Based Interactions

Abstract
This paper presents a novel zone-based flocking control approach suitable for dynamic multi-agent systems (MAS). Inspired by Reynolds behavioral rules for boids, flocking behavioral rules with the zones of repulsion, conflict, attraction, and surveillance are introduced. For each agent, using only bearing and distance measurements, behavioral deviation vectors quantify the deviations from the local separation, local and global flock velocity alignment, local cohesion, obstacle avoidance and boundary conditions, and strategic separation for avoiding alien agents. The control strategy uses the local perception-based behavioral deviation vectors to guide each agent’s motion. Additionally, the control strategy incorporates a directionally-aware obstacle avoidance mechanism that prioritizes obstacles in the agent’s forward path. We also introduce a simplified flocking model where agents interact with their neighbors within a single perception zone, without dividing it into separate interaction zones. Simulation results validate the effectiveness of the models in creating flexible, adaptable, and scalable flocking behavior.
Keywords: Bearing measurement; Multi-agent flocking; Obstacle avoidance
1 Introduction
The concept of flocking draws inspiration from the synchronized collective motion observed in birds, fish, and other social animals. In 1987, Reynolds [1] introduced three rules that govern the flocking behavior of agents, known as boids, in interaction with neighboring agents. These rules, which are as follows, describe the flocking behavior of each agent: $i)$ collision avoidance: avoid collisions with nearby flockmates; $ii)$ velocity matching: attempt to match velocity with nearby flockmates; $iii)$ flock centering: attempt to stay close to nearby flockmates. The original form of these rules is that of verbal expressions, which are general and susceptible to various interpretations. A more mathematical interpretation of these rules is as follows: $i)$ separation: avoid crowding neighbors (within the repulsion zone); $ii)$ alignment: steer towards the average velocities of neighbors (within the attraction zone); $iii)$ cohesion: steer towards the average position of neighbors (within the attraction zone). Flocking studies typically divide the reaction zone of an agent within its perception range into two distinct regions: the zone of repulsion and the zone of attraction [2, 3].
Figure 1: (Left) Flocking behavior based on Reynolds boids model. The current pose of a nominal boid is indicated in solid red, while its future pose, determined by the Reynolds flocking rules, is shown in green. (Middle) Commonly used repulsion and attraction zones model in the literature. (Right) Proposed repulsion, conflict (repulsion and attraction), and attraction zones model.
A three-zone reaction model, which divides the agents’ perception range into non-overlapping zones for repulsion, orientation, and attraction, has also been explored [4]. In these models, agents interact with each other based on the behavioral zone in which their neighbors are located. We introduce a three-zone model to effectively implement the boids behavioral rules with the previously mentioned semi-mathematical interpretation. These zones consist of the zone of repulsion, the zone of conflict (repulsion and attraction), and the zone of attraction. Fig. 1 depicts the behavioral rules in boids within the reaction zone, the conventional two-zone model, and our three-zone model.
Two-zone models typically utilize an attractive–repulsive potential function with a local minimum at a desired separation distance. The potential function exhibits a repulsive force within the repulsion zone and an attractive force within the attraction zone, depending on the pairwise relative distances of neighboring agents [2]. Although the potential field manages both flock aggregation and collision avoidance, it has limitations, including susceptibility to local minima, scalability issues as the number of agents increases, and rigid flock formation.
Olfati-Saber [5] sought to express these behavioral rules mathematically, although not by their direct mathematical translations. Three types of agents, $\alpha$-agents (cooperative agents), $\beta$-agents (obstacles), and $\gamma$-agents (leaders) are introduced. Pairwise potential-based control laws are designed for $\alpha$-agent interactions to form $\alpha$-lattices, typically hexagonal, and to manage interactions with obstacles and leaders. The $\alpha$-lattice flock formation has gained significant attention [6, 7, 8]; however, it lacks a key property of boids flocking behavior—flexible and adaptable spatial configuration. The boids behavioral rules do not dictate a lattice-type rigid spatial formation. The Vicsek [9] and Cucker-Smale [10] models have been broadly acclaimed for exhibiting cohesive flocking and consensus formation. Nevertheless, both models originally lacked flock centering, inter-agent separation, or obstacle avoidance. Moreover, these models led to rigid and occasionally regular patterns, lacking the flexibility of the boids model.
Most existing results aimed at flocking and formation control are based on the assumption that agents exchange their position and velocity information via a high-rate communication channel, typically represented by directed or undirected graphs [11, 12, 13]. While this may be excessive for current multi-agent and multi-robot systems, obstacles and delays impact the communication quality [14]. Furthermore, the limitations and unreliability of global positioning systems drive the need for alternative flocking approaches that do not depend on communication and instead utilize only the agents’ onboard capabilities. On top of that, the original flocking behavioral rules designed for perception-based boids can be more closely replicated.
This paper introduces a novel flocking control approach based on local perception with zone-based interactions. Inspired by the Reynolds rules for realistic flocking behavior, each agent’s perception range is partitioned into the zones of repulsion, conflict, attraction, and surveillance. By introducing the zone of conflict, where both repulsion and attraction coexist, fluid-like motion is enabled. The outer zone of surveillance is to detect and avoid alien agents. The interaction rules define each agent’s behavior, from interactions with nearby flockmates to avoidance of alien agents, obstacles, or boundaries, and adaption to the global flock speed. The interplay of repulsive and attractive forces created by these rules produces the collective motion characteristic of real-world flocking behavior. The desired flocking behavior for each agent is quantified within these zones with simple and perception-based expressions relying solely on bearing and distance measurements. The agent’s control input is directly proportional to its deviations from the desired behaviors. This distributed control strategy allows MAS to achieve flexible and adaptive flocking behavior.
The rest of this paper is organized as follows. Section 2 provides the zone-based flocking model, including the flocking rules, their mathematical translations, and the control strategy. Section 3 derives the flocking control strategy using only bearing and distance measurements. In Section 4, a simplified perception-based flocking model is introduced for the MAS that does not require explicitly defined zones. Section  5 illustrates the simulation results. Section 6 concludes the paper and outlines future work.
2 Zone-Based Flocking Model
2.1 Flocking Rules
For each individual, four concentric zones are defined: the innermost zone of repulsion, the inner annular region of conflict, the intermediate annular region of attraction, and the outermost annular region of surveillance. Without loss of generality, we consider the repulsion zone to be an interval in 1D, a circular area in 2D, and a spherical volume in 3D, all within a distance (or radius) $r_{i}$. The conflict zone is the inner annular region between the repulsion zone and a larger perceptual zone, extending to a distance (or radius) $c_{i}$ from the individual. The attraction zone is the intermediate annular region between the conflict zone and a further perceptual zone, extending to a distance (or radius) $a_{i}$ from the individual. Finally, the surveillance zone is the outermost annular region, extending to a distance (or radius) $s_{i}$ from the individual. Briefly, the reaction zone for each individual is divided into four zones as follows:
  1. 1.
    Repulsion: the innermost zone,
  2. 2.
    Conflict: the inner annular region,
  3. 3.
    Attraction: the intermediate annular region,
  4. 4.
    Surveillance: the outermost annular region.
These zones are illustrated in Fig. 2.
Figure 2: Repulsion, conflict (repulsion and attraction), attraction, and surveillance (conditional repulsion) zones. The set of the neighbors of agent $i$ in the repulsion, conflict, attraction, and surveillance zones, respectively, are $\mathcal{N}_{i}^{r}=\{l\}$, $\mathcal{N}_{i}^{c}=\{s,m\}$, $\mathcal{N}_{i}^{a}=\{f,g\}$, $\mathcal{N}_{i}^{s}=\{o\}$.
In our model, group flocking emerges from the following individual behavioral rules, based on local perception in the zones defined above. These rules are as follows:
  1. 1.
    Local separation: Move away from neighbors in the repulsion and conflict zones until they are positioned at the boundary of the latter zone,
  2. 2.
    Local cohesion: Move to the weighted average position of the neighbors in the conflict and attraction zones,
  3. 3.
    Local alignment: Move to the weighted average velocities of the neighbors in the conflict and attraction zones,
  4. 4.
    Strategic separation: Move away from alien agents until they are positioned at the boundary of the surveillance zone,
  5. 5.
    Obstacle avoidance and boundary constraints: Move away from obstacles and boundaries in the repulsion and conflict zones until they are positioned at the boundary of the latter zone,
  6. 6.
    Global speed alignment: Move to the global desired flock speed.
In the Reynolds boids model, separation is local, with agents maintaining a close yet safe distance from flockmates to ensure cohesion without overcrowding. In addition to local separation, strategic separation involves adopting a larger distance from potentially dangerous alien agents to ensure safety and gain a tactical edge. Local alignment shifts individual speeds toward those of their neighbors, creating a feedback loop that eventually stabilizes the flock’s global speed. When some agents alter their speeds, their neighbors adjust accordingly, resulting in a new global speed. For smooth and coordinated group motion in a flock, it is essential to maintain a cohesive global speed tailored to the flock’s specific purpose, such as foraging or escaping predators. We model natural predator phenomena as alien agents, which indisputably cause collisions, fragmentation, and the dispersal of flock agents. These interaction rules define each agent’s behavior, from interactions with nearby flockmates to avoidance of threats, obstacles, or boundaries, and adaption with the global flock speed. The interplay of repulsive and attractive forces created by these rules produces the collective motion characteristic of real-world flocking behavior.
2.2 Mathematical Expressions of Flocking Rules
Consider a MAS of $n$ agents indexed by $\mathcal{V}=\{1,\cdots,n\}$. Let $\mathbf{p}_{i},\mathbf{v}_{i}\in\mathbb{R}^{d}$ ($d=1,2,3$) denote the position and velocity of the $i$-th agent, respectively, with $i\in\mathcal{V}$. The set of the neighbors of agent $i\in\mathcal{V}$ in its repulsion, conflict, attraction, and surveillance zones are, respectively, defined as $\mathcal{N}_{i}^{r}=\{j|j\neq i\text{ and }\lVert\mathbf{p}_{j}-\mathbf{p}_{i}\rVert\leq r_{i}\}$, $\mathcal{N}_{i}^{c}=\{j|j\neq i\text{ and }r_{i}<\lVert\mathbf{p}_{j}-\mathbf{p}_{i}\rVert\leq c_{i}\}$, $\mathcal{N}_{i}^{a}=\{j|j\neq i\text{ and }c_{i}<\lVert\mathbf{p}_{j}-\mathbf{p}_{i}\rVert\leq a_{i}\}$, $\mathcal{N}_{i}^{s}=\{k|\text{ such that }a_{i}<\lVert\mathbf{q}_{k}-\mathbf{p}_{i}\rVert\leq s_{i}\}$, where $\mathbf{q}_{k}$ denotes the position of $k$-th alien agent and $\lVert$.$\rVert$ . ∥ is the Euclidean norm. Note that $\mathcal{N}_{i}^{r}\cap\mathcal{N}_{i}^{c}\cap\mathcal{N}_{i}^{a}\cap\mathcal{N}_{i}^{s}=\emptyset$. For the sake of brevity, we define $\bar{\mathcal{N}}_{i}^{r}=\mathcal{N}_{i}^{r}\cup\mathcal{N}_{i}^{c}$ and $\bar{\mathcal{N}}_{i}^{a}=\mathcal{N}_{i}^{c}\cup\mathcal{N}_{i}^{a}$.
For directionally aware and flexible collision avoidance with obstacles and boundary constraints, a triangular obstacle avoidance reaction zone is defined around each agent, enclosed within its conflict zone, as illustrated in Fig. 3. The longest altitude of the triangle is aligned with the agent’s heading. Thus, the obstacle avoidance reaction zone is larger in the direction of motion to accommodate the agent’s need for more space to react to obstacles ahead due to its velocity. The opposite side to the longest altitude narrows the reaction zone on the agent’s sides, where collisions with stationary obstacles are less likely. The configuration of the triangle (or tetrahedron in 3D) obstacle avoidance zone can adapt to the agent’s speed or mission. At higher speeds, the longest altitude may extend (e.g., to the agent’s attraction zone) for more reaction space, while at slower speeds, the zone may contract (e.g., to the agent’s repulsion zone).
Figure 3: Triangular obstacle avoidance zone aligned with the agent’s heading.
For each agent, the heading vector is defined as
$\mathbf{h}_{i}=c_{i}\frac{\mathbf{v}_{i}}{\lVert\mathbf{v}_{i}\rVert},$
where $\mathbf{h}_{i}$ has a magnitude of $c_{i}$. For 2D environments, two perpendicular vectors to $\mathbf{h}_{i}$, formed by rotating $\mathbf{h}_{i}$ by ±$\pm\frac{\pi}{2}$, are
$\mathbf{h}_{i}{{}^{\perp}_{+}}=\alpha c_{i}\mathbf{R}(-\frac{\pi}{2})\frac{\mathbf{h}_{i}}{\lVert\mathbf{h}_{i}\rVert},\quad\mathbf{h}_{i}{{}^{\perp}_{-}}=\alpha c_{i}\mathbf{R}(\frac{\pi}{2})\frac{\mathbf{h}_{i}}{\lVert\mathbf{h}_{i}\rVert},$
each with the magnitude of $\alpha c_{i}$ where $0<\alpha\leq 1$ is a scaling factor and $\mathbf{R}$ is the rotation matrix
$\mathbf{R}(\theta)=\begin{bmatrix}\cos{(\theta)}&-\sin{(\theta)}\\ \sin{(\theta)}&\cos{(\theta)}\end{bmatrix}.$
The heading vector $\mathbf{h}_{i}$ represents the longest altitude of the triangular reaction zone. The vectors $\mathbf{h}_{i}^{\perp_{+}}$ and $\mathbf{h}_{i}^{\perp_{-}}$ correspond to the halves of the opposite side to the longest altitude of this zone in 2D. For 3D environments, these halves extend to the vertical dimension, this is crucial for avoiding objects in all directions.
The mathematical translations of the flocking rules are as follows:
  1. 1.
    Local separation:
    $\mathbf{e}_{i}^{ls}=$ $\sum_{j\in\bar{\mathcal{N}}_{i}^{r}}\omega_{ij}^{ls}(\mathbf{p}_{j}-\mathbf{p}_{i}-c_{i}\frac{\mathbf{p}_{j}-\mathbf{p}_{i}}{\lVert\mathbf{p}_{j}-\mathbf{p}_{i}\rVert}),$ (1)
    where $\omega_{ij}^{lr}>0$ is the local separation weight that the agent $i$ assigns to each of its neighbors in $\bar{\mathcal{N}}_{i}^{r}$. The bearing vector $\frac{\mathbf{p}_{j}-\mathbf{p}_{i}}{\lVert\mathbf{p}_{j}-\mathbf{p}_{i}\rVert}$ between agents $i$ and $j$ carries only directional information. When scaled by a desired distance $c_{i}$, it represents the safe or comfortable local separation offset vector for agent $i$ relative to neighboring agent $j$ within agent $i$’s repulsion and conflict zones. The vector $\mathbf{e}_{i}^{ls}\in\mathbb{R}^{d}$ denotes agent $i$’s deviations from the desired local separation behavior and creates a repulsive potential for that agent.
  2. 2.
    Local cohesion:
    $\mathbf{e}_{i}^{lc}=\frac{1}{\sum_{j\in\bar{\mathcal{N}}_{i}^{a}}\omega_{ij}^{lc}}\sum_{j\in\bar{\mathcal{N}}_{i}^{a}}\omega_{ij}^{lc}\mathbf{p}_{j}-\mathbf{p}_{i},$ (2)
    where $\omega_{ij}^{lc}>0$ is the local cohesion weight that the agent $i$ assigns to each of its neighbors in $\bar{\mathcal{N}}_{i}^{a}$. The vector $\mathbf{e}_{i}^{lc}\in\mathbb{R}^{d}$ denotes agent $i$’s vector deviations from the desired local cohesion behavior and creates an attractive potential for that agent.
  3. 3.
    Local alignment:
    $\mathbf{e}_{i}^{la}=\frac{1}{\sum_{j\in\bar{\mathcal{N}}_{i}^{a}}\omega_{ij}^{la}}\sum_{j\in\bar{\mathcal{N}}_{i}^{a}}\omega_{ij}^{la}\mathbf{v}_{j}-\mathbf{v}_{i},$ (3)
    where $\omega_{ij}^{la}>0$ is the local alignment weight that the agent $i$ assigns to each of its neighbors in $\bar{\mathcal{N}}_{i}^{a}$. The vector $\mathbf{e}_{i}^{la}\in\mathbb{R}^{d}$ denotes agent $i$’s deviations from the desired local alignment behavior and creates an attractive potential.
  4. 4.
    Strategic separation:
    $\mathbf{e}_{i}^{ss}$ $=\sum_{k\in\mathcal{N}_{i}^{s}}\omega_{ik}^{ss}(\mathbf{q}_{k}-\mathbf{p}_{i}-s_{i}\frac{\mathbf{q}_{k}-\mathbf{p}_{i}}{\lVert\mathbf{q}_{k}-\mathbf{p}_{i}\rVert}),$ (4)
    where $\omega_{ik}^{ss}>0$ is the strategic separation weight that the agent $i$ assigns to each alien agent (i.e., non-flockmate agent) in $\mathcal{N}_{i}^{s}$. The strategic separation offset vector $s_{i}\frac{\mathbf{q}_{j}-\mathbf{p}_{i}}{\lVert\mathbf{q}_{j}-\mathbf{p}_{i}\rVert}$ represents a safe directional distance from a potentially dangerous alien $k$. The vector $\mathbf{e}_{i}^{ss}\in\mathbb{R}^{d}$ denotes agent $i$’s deviations from the desired strategic separation behavior and creates a repulsive potential.
  5. 5.
    Obstacle avoidance and boundary constraints:
    $\mathbf{e}_{i}^{oa}$ $=\omega_{i}^{oa}(\mathbf{b}-\mathbf{p}_{i}-c_{i}\frac{\mathbf{b}-\mathbf{p}_{i}}{\lVert\mathbf{b}-\mathbf{p}_{i}\rVert}),$ (5)
    where $\omega_{i}^{oa}>0$ is the obstacle avoidance weight that the agent $i$ assigns to the obstacle within its obstacle avoidance reaction zone and $\mathbf{b}$ denotes the closest point on the obstacle’s edge or boundary to the agent. The obstacle avoidance offset vector $c_{i}\frac{\mathbf{b}-\mathbf{p}_{i}}{\lVert\mathbf{b}-\mathbf{p}_{i}\rVert}$ represents a safe directional distance from obstacles and boundaries. The vector $\mathbf{e}_{i}^{oa}\in\mathbb{R}^{d}$ denotes agent $i$’s deviations from the desired obstacle avoidance behavior and creates a repulsive potential.
    Each agent continuously monitors whether its obstacle avoidance reaction zone intersects with an obstacle or boundary. A straightforward approach is to check if its heading vector $\mathbf{h}_{i}$ and the perpendicular vectors (i.e., in 2D, the vectors $\mathbf{h}_{i}{{}^{\perp}_{+}}$ and $\mathbf{h}_{i}{{}^{\perp}_{-}}$) intersect with an obstacle or boundary. If an intersection is detected, the agent identifies $\mathbf{b}$. The offset vector $\mathbf{e}_{i}^{oa}$ is then computed proportionally to the distance from the boundary and is used to adjust the agent’s position accordingly. The triangular zone-based directional obstacle avoidance strategy focuses on obstacles directly ahead in the agent’s trajectory. By prioritizing avoidance in the heading direction, the agent minimizes unnecessary reactions to obstacles that are less likely to impact its path.
  6. 6.
    Global speed alignment:
    $\mathbf{e}_{i}^{ga}=\omega_{i}^{ga}(v^{d}\frac{\mathbf{v}_{i}}{\lVert\mathbf{v}_{i}\rVert}-\mathbf{v}_{i}),$ (6)
    where $v_{d}$ denotes the desired global speed and $\omega_{i}^{ga}\geq 0$ is the global speed alignment weight or gain. If agent $i$ is informed about $v_{d}$, then $\omega_{i}^{ga}>0$, otherwise $\omega_{i}^{ga}=0$. For each informed agent $i$, the vector $\mathbf{e}_{i}^{ga}\in\mathbb{R}^{d}$ denotes its deviations from the desired global speed alignment behavior and creates an attractive potential.
Remark 1.
If the agent’s conflict and surveillance zones are not defined, $c_{i}$ and $s_{i}$ default to $r_{i}$ in (1),(4), and (5).
Remark 2.
Given the magnitude of the position vectors, it is imperative to scale down the magnitude of $\mathbf{e}_{ij}^{lc}$ while preserving its direction to ensure consistency. One straightforward method is to normalize $\mathbf{e}_{i}^{lc}$, and then multiply it by a positive weight.
2.3 Flocking Controller
The deviation of an agent from the desired flocking behavior, as defined in (1)–(6), represents the flocking requirements for that agent. The total deviation vector for agent $i\in\mathcal{V}$ from the desired flocking behavior is expressed as
$\mathbf{e}_{i}=\mathbf{e}_{i}^{ls}+\mathbf{e}_{i}^{lc}+\mathbf{e}_{i}^{la}+\mathbf{e}_{i}^{ss}+\mathbf{e}_{i}^{oa}+\mathbf{e}_{i}^{ga}.$ (7)
Let $\mathbf{u}_{i}\in\mathit{\mathbb{R}}^{d}$ denote the control input vector for agent $i\in\mathcal{V}$. We define
$\mathbf{u}_{i}=g_{i}\mathbf{e}_{i},$ (8)
where $g_{i}$ is the controller gain. Unifying each agent’s repulsion and attraction potentials into the control input creates an artificial flocking potential energy for that agent.
Remark 3.
With boundary conditions, the environment is confined, and thus, even if some agents become separated from the flock or the flock fragments, the agents will eventually reunite. In the absence of boundary conditions, it is necessary to assume that each agent always has at least one flockmate within its conflict or attraction zone. To ensure this, a straightforward approach is to gradually increase the radius of the attraction zone until a flockmate is detected.
Remark 4.
The controller (8) does not explicitly account for bounded control effort. The control input of agent $i$ are bounded as follows,
$\mathbf{u}_{i}=u_{i}^{\max}\tanh{\left(\frac{\lVert\mathbf{u}_{i}\rVert}{u_{i}^{\max}}\right)}\frac{\mathbf{u}_{i}}{\lVert\mathbf{u}_{i}\rVert},$ (9)
where $u_{i}^{\max}$ are the maximum acceleration magnitudes for agent $i$. These smooth cutoffs at $u_{i}^{\max}$ maintain the direction while reducing the magnitude.
Remark 5.
The resultant velocity from (9) for physical agents is subject to saturation. To prevent velocity saturation, the velocity of agent $i$ is restricted as follows,
$\mathbf{v}_{i}=v_{i}^{\max}\tanh{\left(\frac{\lVert\mathbf{v}_{i}\rVert}{v_{i}^{\max}}\right)}\frac{\mathbf{v}_{i}}{\lVert\mathbf{v}_{i}\rVert},$ (10)
where $v_{i}^{\max}$ is the maximum velocity magnitudes for agent $i$. If the computed velocity exceeds the agent’s limits, it is smoothly capped at $v_{i}^{\max}$.
3 Flocking with Bearing-Distance Measurements
In general, MAS control can be classified into position-based, displacement-based, and distance-based control [15]. The controller in (8) falls under the position-based category. In this approach, agents sense their positions relative to a global coordinate system and actively adjust their positions to meet the desired behavior, defined in terms of desired positions within the global coordinate system. Consequently, this method requires agents to be equipped with precise global positioning sensors. Let $\mathbf{p}_{ij}=\mathbf{p}_{j}-\mathbf{p}_{i}$ be the relative position vector , with $\dot{\mathbf{p}}_{ij}=\mathbf{v}_{j}-\mathbf{v}_{i}$. If the relative positions $\mathbf{p}_{ij}$ are sensed with respect to the global coordinate system, meaning the agents know the orientation of the global coordinate system, all flocking behavioral deviation vectors can be re-expressed in terms of $\mathbf{p}_{ij}$ and $\dot{\mathbf{p}}_{ij}$. As a result, the controller in (8) transitions to a displacement-based control approach, where agents no longer need to sense their absolute positions. In distance-based control, agents actively regulate the distances between themselves and their neighbors to achieve the desired behavior defined by the inter-agent distances. Each agent senses the relative positions of its neighbors with respect to its own local coordinate system, with the orientations of these local coordinate systems not necessarily being aligned. While distance-based control offers advantages in terms of reduced sensing requirements, it demands more interactions among agents to maintain the desired behavior. Since flocking agents, like boids, are perception-driven interactive entities, displacement-based and distance-based control approaches are more practical for flocking. In the following, we derive the desired behavioral deviation vectors for flocking based on bearing measurements and inter-agent distances. This eliminates the need for agents to have knowledge of the global coordinate system or their positions.
Let the bearing vector $\mathbf{b}_{ij}=\frac{\mathbf{p}_{j}-\mathbf{p}_{i}}{\lVert\mathbf{p}_{j}-\mathbf{p}_{i}\rVert}$ and inter-agent distance $d_{ij}=\lVert\mathbf{p}_{j}-\mathbf{p}_{i}\rVert$ be available from the onboard sensors. Their rate of change is also either measured directly or computed from the sensor data. The bearing vector $\mathbf{b}_{ij}$ and inter-agent distance $d_{ij}$ are related by
$\mathbf{b}_{ij}=\frac{\mathbf{p}_{j}-\mathbf{p}_{i}}{d_{ij}}.$ (11)
The time derivative of the bearing vector $\mathbf{b}_{ij}$ is given by
$\dot{\mathbf{b}}_{ij}=\frac{1}{d_{ij}}(\mathbf{v}_{j}-\mathbf{v}_{i})-\dot{d}_{ij}\mathbf{b}_{ij},$
or,
$\mathbf{v}_{j}-\mathbf{v}_{i}=d_{ij}(\dot{\mathbf{b}}_{ij}+\dot{d}_{ij}\mathbf{b}_{ij}).$ (12)
The desired behavioral deviation vectors for flocking, based on bearing measurements and inter-agent distances, are derived as follows:
  1. 1.
    Bearing-distance based local separation: By substituting the bearing vector $\mathbf{b}_{ij}$ into (1) and using (11), we obtain the bearing-distance based local separation deviation vector for the agent $i$ as
    $\mathbf{e}_{i}^{ls}=$ $\sum_{j\in\bar{\mathcal{N}}_{i}^{r}}\omega_{ij}^{ls}(d_{ij}-c_{i})\mathbf{b}_{ij}.$ (13)
  2. 2.
    Bearing-distance based local cohesion: Multiplying both sides of local cohesion in (2) by $\sum_{j\in\bar{\mathcal{N}}_{i}^{la}}\omega_{ij}^{lc}$ yields
    $\sum_{j\in\bar{\mathcal{N}}_{i}^{a}}\omega_{ij}^{lc}\mathbf{e}_{i}^{lc}$ $=\sum_{j\in\bar{\mathcal{N}}_{i}^{a}}\omega_{ij}^{lc}\mathbf{p}_{j}-\sum_{j\in\bar{\mathcal{N}}_{i}^{a}}\omega_{ij}^{lc}\mathbf{p}_{i}$
    $=\sum_{j\in\bar{\mathcal{N}}_{i}^{a}}\omega_{ij}^{lc}(\mathbf{p}_{j}-\mathbf{p}_{i})=\sum_{j\in\bar{\mathcal{N}}_{i}^{a}}\omega_{ij}^{lc}d_{ij}\mathbf{b}_{ij},$
    or,
    $\mathbf{e}_{i}^{lc}=\frac{1}{\sum_{j\in\bar{\mathcal{N}}_{i}^{a}}\omega_{ij}^{lc}}\sum_{j\in\bar{\mathcal{N}}_{i}^{la}}\omega_{ij}^{lc}d_{ij}\mathbf{b}_{ij},$ (14)
    which is the bearing-distance based local cohesion deviation vector for the agent $i$.
  3. 3.
    Bearing-distance based local alignment: Similarly to the bearing-distance based local cohesion, we arrive at
    $\mathbf{e}_{i}^{la}=\frac{1}{\sum_{j\in\bar{\mathcal{N}}_{i}^{a}}\omega_{ij}^{la}}\sum_{j\in\bar{\mathcal{N}}_{i}^{a}}\omega_{ij}^{la}(\mathbf{v}_{j}-\mathbf{v}_{i}).$
    Using (12), the bearing-distance based local alignment deviation vector for agent $i$ is given by
    $\mathbf{e}_{i}^{la}=\frac{1}{\sum_{j\in\bar{\mathcal{N}}_{i}^{a}}\omega_{ij}^{la}}\sum_{j\in\bar{\mathcal{N}}_{i}^{a}}\omega_{ij}^{la}d_{ij}(\dot{\mathbf{b}}_{ij}+\dot{d}_{ij}\mathbf{b}_{ij}).$ (15)
  4. 4.
    Bearing-distance based strategic separation: Denote $\mathbf{b}_{ik}$ as the bearing vector between agent $i$ and an alien agent $k$, and $d_{ik}$ as the distance between them. The bearing-distance based strategic separation deviation vector for agent $i$ is given by
    $\mathbf{e}_{i}^{ss}=$ $\sum_{j\in\mathcal{N}_{i}^{s}}\omega_{ik}^{ss}(d_{ik}-s_{i})\mathbf{b}_{ik}.$ (16)
  5. 5.
    Bearing-distance based obstacle avoidance and boundary constraints: Denote $\mathbf{b}_{ie}$ as the bearing vector between agent $i$ and a boundary edge (of an obstacle or environment), and $d_{ie}$ as the distance between them. The bearing-distance based obstacle avoidance and boundary constraints deviation vector for agent $i$ is given by
    $\mathbf{e}_{i}^{oa}$ $=\omega_{i}^{oa}(d_{ie}-c_{i})\mathbf{b}_{ie}.$ (17)
  6. 6.
    Bearing-distance based global speed alignment: Define the auxiliary vector
    $\mathbf{u}_{i}^{\text{aux}}=g_{i}(\mathbf{e}_{i}^{ls}+\mathbf{e}_{i}^{lc}+\mathbf{e}_{i}^{la}+\mathbf{e}_{i}^{ss}+\mathbf{e}_{i}^{oa}).$ (18)
    For an informed agent $i$ with the desired global speed $v^{d}$, its velocity is estimated based on the bearing-distance measurements as follows,
    $\hat{\mathbf{v}}_{i}=\hat{\mathbf{v}}_{i}^{\text{prev}}+\delta\mathbf{u}_{i}^{\text{aux}},$ (19)
    where $\hat{\mathbf{v}}_{i}^{\text{prev}}$ denotes the previously estimated velocity at the prior time step, and $\delta$ is the time step duration. Accordingly, the bearing-distance based global speed alignment deviation vector for agent $i$ is given by
    $\mathbf{e}_{i}^{ga}=\omega_{i}^{ga}(v^{d}\frac{\hat{\mathbf{v}}_{i}}{\lVert\hat{\mathbf{v}}_{i}\rVert}-\hat{\mathbf{v}}_{i}).$ (20)
Given that all behavioral deviation vectors are derived in terms of bearing-distance measurements, the bearing-distance based flocking control input vector for agent $i\in\mathcal{V}$ is computed as
$\mathbf{u}_{i}=\mathbf{u}_{i}^{\text{aux}}+g_{i}\mathbf{e}_{i}^{ga}.$ (21)
The proposed distributed flocking scheme is summarized in Algorithm 1. For each agent, the behavioral deviation vectors for local separation, local cohesion, local alignment, strategic separation, and obstacle avoidance using bearing vectors and inter-agent distances. These deviation vectors are aggregated to form an auxiliary vector to estimate the agent’s velocity in alignment with a desired global speed. The control input for each agent is then computed by combining the auxiliary vector with the global speed alignment vector. This distributed approach enables adaptive flocking control based on local interactions. In the algorithm, $m$ denotes the number of alien agents, $o$ represents the number of obstacles, and $\mathbf{p}^{\min}$ and $\mathbf{p}^{\max}$ specify the boundary conditions.
1:  Set $d$, $n$, $v_{d}$, $m$, $o$, $\mathbf{p}^{\min},\mathbf{p}^{\max}$.
2:  Initialize $o$ obstacles.
3:  Initialize $\mathbf{p}_{i}$ for $i\in\{1,\cdots,n\}$ and $\mathbf{p}_{k}^{a}$ for $k\in\{1,\cdots,m\}$.
4:  Initialize $r_{i}$, $c_{i}$, $a_{i}$, $s_{i}$, $g_{i}$, $v_{i}^{\max}$, $\mathbf{v}_{i}$, $u_{i}^{\max}$, $\omega_{ij}^{ls}$, $\omega_{ij}^{lc}$, $\omega_{ij}^{la}$, $\omega_{ik}^{ss}$, $\omega_{i}^{oa}$, $\omega_{i}^{ga}$ for $i,j\in\{1,\cdots,n\}$, and $v_{k}^{a\max}$, $\mathbf{v}_{k}^{a}$ for $k\in\{1,\cdots,m\}$.
5:  for each time step do
6:     for each agent $i\in\{1,\cdots,n\}$ do
7:        Identify neighbors $\mathcal{N}_{i}^{r}$, $\mathcal{N}_{i}^{c}$, $\mathcal{N}_{i}^{a}$, and $\mathcal{N}_{i}^{s}$ based on distance measurements $d_{ij}$, $d_{ik}$, $d_{ie}$.
8:        Measure the bearing vectors $\mathbf{b}_{ij}$, $\mathbf{b}_{ik}$, $\mathbf{b}_{ie}$.
9:        Calculate separation deviation $\mathbf{e}_{i}^{ls}$ from (13).
10:        Calculate cohesion deviation $\mathbf{e}_{i}^{lc}$ from(14).
11:        Calculate alignment deviation $\mathbf{e}_{i}^{la}$ from (15).
12:        Calculate strategic separation deviation $\mathbf{e}_{i}^{ss}$ from (16).
13:        Calculate obstacle avoidance and boundary conditions deviation $\mathbf{e}_{i}^{oa}$ in (17).
14:        Calculate auxiliary $\mathbf{u}_{i}^{\text{aux}}$ from (18).
15:        Calculate $\hat{\mathbf{v}}_{i}$ from (19).
16:        Calculate global speed alignment deviation $\mathbf{e}_{i}^{ga}$ from (20).
17:        Calculate control input $\mathbf{u}_{i}$ from (21).
18:        Bound control input $\mathbf{u}_{i}$ from (9).
19:        Update velocity $\mathbf{v}_{i}$.
20:        Bound velocity $\mathbf{v}_{i}$ from (10).
21:        Update position $\mathbf{p}_{i}$.
22:     end for
23:     for each alien $k\in\{1,\cdots,m\}$ do
24:        Update $\mathbf{p}_{k}^{a}$.
25:     end for
26:  end for
Algorithm 1 Bearing-Distance Based Flocking with Zone-Based Interactions
Boids-like flocking is unstructured, allowing agents to rearrange freely as long as they maintain cohesion, separation, and alignment. In contrast, lattice-type flocking enforces a regular geometric arrangement.
Remark 6.
When $c_{i}=c\forall i\in\mathcal{V}$, and incorporating the separation (13) alongside the alignment (15) while omitting the cohesion (14), the flock adopts a lattice-type geometric configuration.
4 Simplified Perception-Based Flocking
In certain MAS or scenarios, the previously proposed flocking with zone-based interactions might be excessive, impractical, or computationally inefficient. In the following, we introduce a simplified flocking model in which the agents interact with their neighbors within their perception zone, without partitioning it into distinct interaction zones. The simplified perception-based flocking rules are as follows:
  1. 1.
    separation–cohesion:
    $\mathbf{e}_{i}^{sc}=\sum_{j\in\mathcal{N}_{i}}\omega_{ij}^{sc}(\mathbf{p}_{j}-\mathbf{p}_{i}-\alpha_{i}r_{i}\frac{\mathbf{p}_{j}-\mathbf{p}_{i}}{\lVert\mathbf{p}_{j}-\mathbf{p}_{i}\rVert})=\sum_{j\in\mathcal{N}_{i}}\omega_{ij}^{sc}(d_{ij}-\alpha_{i}r_{i})\mathbf{b}_{ij},$ (22)
  2. 2.
    alignment:
    $\mathbf{e}_{i}^{a}=\frac{1}{\sum_{j\in\mathcal{N}_{i}}\omega_{ij}^{a}}\sum_{j\in\mathcal{N}_{i}}\omega_{ij}^{a}\mathbf{v}_{j}-\mathbf{v}_{i}=\frac{1}{\sum_{j\in\mathcal{N}_{i}}\omega_{ij}^{a}}\sum_{j\in\mathcal{N}_{i}}\omega_{ij}^{a}d_{ij}(\dot{\mathbf{b}}_{ij}+\dot{d}_{ij}\mathbf{b}_{ij}),$ (23)
where $\alpha_{i}=\beta_{i}|\mathcal{N}_{i}|$, $\beta_{i}$ is a scaling factor, typically chosen to be inversely proportional to $r_{i}$, e.g., $\beta_{i}=\frac{1}{r_{i}}$, $r_{i}$ is the radius of the perception zone for agent $i$, $\mathcal{N}_{i}$ denotes the set of neighbors of agent $i$ within its perception zone, $|\mathcal{N}_{i}|$ denotes the cardinality of $\mathcal{N}_{i}$, and $\omega_{ij}^{sc},\omega_{ij}^{a}>0$. The simplified flocking controller is given by
$\mathbf{u}_{i}=g_{i}(\mathbf{e}_{i}^{sc}+\mathbf{e}_{i}^{a}).$
The separation–cohesion rule in (22) alone results in oscillatory behavior between separation and cohesion; agents continually move away from neighbors within the range $\alpha_{i}r_{i}$ and then approach them once they are outside this range. However, this rule by itself does not help agents establish a common motion direction or stabilize a sufficient separation distance to prevent collisions. Incorporating the alignment rule in (23) addresses both issues, enabling agents to achieve coordinated motion and avoid close encounters.
In a homogeneous MAS, we can set $\omega_{ij}^{sc}=\omega_{ij}^{a}=g_{i}=1$ for all $i,j\in\{1,\cdots,n\}$. Consequently, homogeneous MAS flocking emerges with either the position-based controller
$\mathbf{u}_{i}=\sum_{j\in\mathcal{N}_{i}}(\mathbf{p}_{j}-\mathbf{p}_{i}-\alpha_{i}r_{i}\frac{\mathbf{p}_{j}-\mathbf{p}_{i}}{\lVert\mathbf{p}_{j}-\mathbf{p}_{i}\rVert})+\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{v}_{j}-\mathbf{v}_{i},$ (24)
or the bearing-distance-based controller
$\mathbf{u}_{i}=\sum_{j\in\mathcal{N}_{i}}(d_{ij}-\alpha_{i}r_{i})\mathbf{b}_{ij}+\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}d_{ij}(\dot{\mathbf{b}}_{ij}+\dot{d}_{ij}\mathbf{b}_{ij}).$ (25)
We can derive the position-based and bearing-distance-based controllers above from each other, rendering them equivalent. Therefore, in scenarios where bearing-distance measurements are unavailable but position data is accessible, the position-based controller can achieve flocking behavior consistent with that of the bearing-distance-based controller.
Remark 7.
In complex environments, the simplified perception-based model can incorporate strategic separation, obstacle avoidance and boundary constraints, global speed alignment, and even local cohesion by adapting each to the single perception zone-based flocking scenario.
4.1 Advantages of the Proposed Flocking Model
Traditional flocking models emphasize different aspects of flocking behavior, such as local attraction-repulsion in the Olfati-Saber model [5], simple directional alignment in the Vicsek model [9], and distance-dependent velocity averaging in the Cucker-Smale model [10]. Our simplified flocking model (e.g., the position-based model in (24) for comparison), innovatively integrates all three emergent flocking behaviors: cohesion-separation (corresponding to attraction-repulsion) and alignment (neighborhood velocity averaging). Our approach offers several key advantages. In comparison to:
  1. 1.
    Olfati-Saber Model: $i$) (Direct Cohesion-Separation) Unlike Olfati-Saber’s indirect approach using a relative distance-based potential function, our model directly and straightforwardly enforces cohesion and separation. $ii$) Alignment Definition: While Olfati-Saber emphasizes velocity consensus, our model focuses on average velocities. $iii$) Flexible Spatial Arrangement: Flexible Spatial Arrangement: The Olfati-Saber model typically produces lattice-type flock formations—likely due to the influence of the velocity consensus term—which limits flexibility in spatial arrangement. In contrast, our model achieves the adaptable configurations seen in biological flocks by incorporating a neighborhood velocity averaging term.
  2. 2.
    Vicsek Model: This model assumes constant speed and aligns velocity direction without explicitly incorporating cohesion and separation and varying speeds of individuals. In contrast, our model seamlessly integrates these elements, fostering more natural flocking behavior.
  3. 3.
    Cucker-Smale Model: This model focuses on velocity alignment influenced by relative positions, where alignment strength is distance-dependent, lacking direct enforcing of cohesion and separation. Our model, however, incorporates and balances cohesion (attraction), separation (repulsion), and alignment more effectively, providing a more comprehensive approach to flocking dynamics.
Additionally, our model features a dynamic scaling factor, $\alpha_{i}$, which adapts the separation based on the influence of neighboring agents in relation to their current density. This allows for more flexible interactions in varying crowd conditions, enhancing stability and reducing oscillations in dense environments. By scaling the influence based on the number of neighbors, agents are less likely to be overwhelmed by excessive influences, resulting in more stable flocking behavior. The ability to dynamically adjust interactions based on neighbor density makes our model particularly scalable for real-time applications, especially in robotics and simulations where agent populations may fluctuate. Overall, these advantages position our model as a more effective and adaptable solution for simulating and controlling flocking behavior across a variety of scenarios compared to traditional models.
5 Simulation Results
This section presents several illustrative examples to validate the proposed flocking models introduced in Section 3 and Section 4. Beginning with the model from Section 3—designed for highly complex environments—we consider a MAS with $n=10$ agents, indexed by $\mathcal{V}=\{1,\cdots,10\}$, operating in a 2D environment ($d=2$). Each agent follows the double-integrator dynamics [16]
$\begin{cases}\dot{\mathbf{p}}_{i}=\mathbf{v}_{i},\\ \dot{\mathbf{v}}_{i}=\mathbf{u}_{i},\end{cases}$ (26)
where $\mathbf{u}_{i}\in\mathbb{R}^{2}$ denotes the control input (acceleration) for agent $i\in\mathcal{V}$.
For each agent $i\in\mathcal{V}$, the zone radii are set as follows, the repulsion zone radius $r_{i}=1.5m$, the conflict zone radius $c_{i}=3.0m$, the attraction zone radius $a_{i}=5.0m$, and the surveillance zone radius $s_{i}=6.0m$. The control gains are uniformly set to $g_{i}=1.0$ for all agents. We set weights as follows, repulsion weights $\omega_{ij}^{ls}=5$, cohesion attraction weights $\omega_{ij}^{lc}=0.75$, and alignment attraction weights $\omega_{ij}^{la}=0.25$. Additionally, the surveillance zone repulsion weight $\omega_{ik}^{ss}$ and obstacle avoidance weight $\omega_{i}^{oa}$ are both set to 5. The global attraction weights for each agent are defined as $\omega_{i}^{ga}=5$. The repulsion weights are slightly bigger than the attraction weight to prioritize separation and collision avoidance over cohesion and alignment. The environment is confined, with boundaries from $[0m,0m]$ to $[50m,50m]$. The environment also contains an irregularly shaped stationary obstacle and contains an irregularly shaped stationary obstacle. An alien agent is trapped within a triangular zone within the environment. The flock agents are unaware of the alien’s presence until it enters their surveillance zone. The alien can detect flock agents within a radius of $9m$. Once the flock agents come within range, the alien follows the closest agent until it reaches the borders of its triangular zone. The shorter surveillance radius of the flock agents causes a delayed reaction to the alien’s presence. The agents’ positions and velocities were initialized randomly within the range $(0m,10m)$ and $(0m/s,1m/s)$. The maximum speed and acceleration for each agent are set to $v_{i}^{\text{max}}=4m/s$ and $u_{i}^{\text{max}}=2m/s^{2}$, respectively. The desired global flock speed is defined as $v_{d}=3m/s$. The total time is $100.0s$, with a time step of $0.1s$.
Fig. 4 illustrates the resultant flocking behavior. The alien follows the nearest agent when close, causing temporary fragmentation in the cohesive flock. Agents navigate a narrow passage between the obstacle and boundaries, a maneuver infeasible with conventional circular buffer zone methods.
Figure 4: Flocking behavior of $10$ agents. The flocking environment is confined and includes an irregular polygon-shaped stationary obstacle and an alien agent. The alien (shown in black), constrained within the transparent blue zone, begins to follow the nearest flocking agent when within proximity. The flock initially demonstrates cohesive and collision-free motion; however, the arrival of the alien agent causes a temporary fragmentation, though the flock eventually reunites. The agents navigate through a narrow passage between the stationary obstacle and the environmental boundaries. Such a maneuver would not be permitted with a conventional obstacle avoidance method, which employs a circular buffer zone around irregular obstacles (depicted in gray).
Figure 5 shows the profiles of inter-agent distances, speeds, and control inputs. Inter-agent distances stay above the repulsion radius of $r_{i}=1.5m$, reaching up to 10 meters when cohesive. Fragmentation due to the alien’s interference temporarily increases distance ranges before normalizing. Agent speeds remain below the maximum of $4m/s$, with the average flock speed close to the desired $3m/s$. Control inputs are within the specified bounds of $2m/s^{2}$.
[…]
The simulation example illustrated the emergence of flocking behavior in accordance with the zone-based flocking rules defined in subsection 2.1. The agents, initially positioned and moving randomly, formed a cohesive group, maintaining relative distances while moving as a collective entity. The flock effectively navigated around an irregular obstacle with the proposed directionally aware obstacle avoidance mechanism, which allowed closer proximity to the obstacle’s edges and passage through a narrow gap, maneuvers not feasible with conventional circular buffer zone methods. Despite a temporary fragmentation caused by an alien agent, the flock reformed cohesively. These results highlight the effectiveness of the bearing-distance based flocking approach with zone-based interactions for realistic boids-like flocking.
We also evaluate the simplified perception-based flocking model introduced in Section 4. This assessment includes both 2D and 3D flocking scenarios, each involving $n=60$ homogeneous agents employing the simplified perception-based flocking controller in (25). For each agent $i$, the radius of perception is defined as $r_{i}=10m$. The agents’ positions and velocities were initialized randomly within the range $(0m,30m)$ and $[-1m/s,1m/s]$. The scaling parameters for both 2D and 3D scenarios are set to $\beta_{i}=\frac{1}{r_{i}}$. The total simulation time is $20.0s$ for the 2D case and $10.0s$ for the 3D case, with a time step of $0.1s$. The simulation results, presented in Fig. 6, indicate that the agents successfully achieve cohesive and collision-free flocking behavior. This is evidenced by stabilized inter-agent distances that ensure sufficient separation to avoid collisions. Additionally, the speed profiles demonstrate that the agents align effectively. These results emphasize the simplified perception-based flocking model’s effectiveness in establishing and maintaining a cohesive and aligned flock.
Figure 6: Flocking behavior of $60$ agents for both 2D (top row) and 3D (bottom row) scenarios using the simplified perception-based flocking controller. The cohesive and collision-free flocking behavior is successfully established. Inter-agent distances stabilize while maintaining sufficiently safe separation. The speed profiles of the flock agents indicate achieving alignment.
6 Conclusions
We introduced a distributed perception-based flocking control model that leverages zone-based interactions to achieve adaptable and dynamic flocking behavior without relying on communication or global positioning. By incorporating a conflict zone where repulsive and attractive forces interact, our model captures the fluidity and adaptability of natural flocking. Simulation results validated the model’s ability to create and manage a cohesive flock in complex scenarios, such as navigating confined spaces and interacting with alien agents. Toward less predictable and more lifelike motions like in realistic flocking, stochastic elements, such as stochastic weights or random perturbations in velocities, can be integrated into the model. The model’s reliance on local sensing without communication highlights its potential for scalable applications in real-world autonomous systems.